cs161: Data Structures and Algorithms Handout #15 Michael Goldwasser Stanford University Tuesday, August 6, 1996

Homework #6: Graphs, Spanning Trees, and Shortest Paths Due Date: Tuesday, 13 August 1996

Reading

Read Ch. 24, Ch. 25.

Practice

Recall, these exercises are purely for your own practice. You need only turn in the official problems.

• Do CLR 24.1-5.

• Do CLR 24.1-3, 24.1-4.

• Do CLR 25.2-4.

• Do CLR 25.3-3.

Problems

1. Do CLR 24.2-4 and 24.2-5.

2. Do CLR 25-2.

3. Do CLR 25-4.

Extra Credit

Do CLR 24.2-8.